OpenStudy (anonymous):

Consider the reaction of aluminum chloride with silver acetate. How ml of 0.250M aluminum chloride are required to react completely with 20ml of 0.500M silver acetate?

OpenStudy (anonymous):

AlCl3 + 3(AgC2H3O2) -> Al(C2H3O2)3 + 3(AgCl)

OpenStudy (australopithecus):

Yes I was about to tell you to post that

OpenStudy (australopithecus):

First off recognize that Molarity = moles/Liters

OpenStudy (australopithecus):

You have a volume 20ml of 0.500M acetate

OpenStudy (australopithecus):

Thus you can say x moles/0.02L = 0.500M of acetate solve for the moles

OpenStudy (australopithecus):

Look at the ratio you have 3 times the moles of silver acetate for everyone one mole of aluminum chloride

OpenStudy (australopithecus):

Since the division cancels multiplication you simply divide the number of moles of silver acetate by 3 to solve for the moles of aluminum chloride

OpenStudy (australopithecus):

This will give you the moles, since you have the concentration of aluminum chloride you can just sub them into the equation to solve for the liters Molarity = moles/liters

OpenStudy (australopithecus):

remember the conversion Liters * 1000 = mL mL/1000 = Liters

OpenStudy (australopithecus):

if you have questions about the basic algebra feel free to ask

OpenStudy (anonymous):

Thanks a lot! i'll try this now

OpenStudy (australopithecus):

alright if you have problems show your work and I can help have lecture now though so yeah

OpenStudy (anonymous):

i did the this: 1 mole of AlCl3 -> 3 moles of AgC2H3O2 x -> 0.01 so x= 0.0033. 0.5 = 0.003/volume , volume = 0.0066, is it right?

OpenStudy (australopithecus):

0.500M = x/0.020L x = 0.500M*0.020L x=0.01moles 0.01moles/3 = 0.0033mol of Ag acetate thus 0.250M = 0.0033mol/x 0.0033mol/0.250 = x x = 0.013L

OpenStudy (australopithecus):

0.013L = 13mL

OpenStudy (anonymous):

oh! that it! now i realize that. Thanks !!

OpenStudy (australopithecus):

I think you made a mistake with your algebra remember Molarity = Moles/Volume thus you have 0.250M = 0.0033mol/x = 0.250M*x = 0.0033mol*x/x = 0.250M*x = 0.0033mol = 0.250M*x/0.250M = 0.0033mol/0.250M = x = 0.0033mol/0.250M

OpenStudy (australopithecus):

here are all the steps

OpenStudy (australopithecus):

remember you want to bring the x into the numerator and isolate it

OpenStudy (australopithecus):

you never want it in the denominator